OpenStudy (anonymous):

Help Solve the equation for k. Hint. First Multiply both sides by 2

OpenStudy (anonymous):

\[\frac{ 2mk +2j}{ 2 } = n+p \]

OpenStudy (anonymous):

If you multiplied both sides you would get \[ mk+mj = n+ p\] right?

jimthompson5910 (jim_thompson5910):

actually, it would be \[\Large 2mk+2j = 2(n+p)\]

OpenStudy (anonymous):

ok 2mk+ 2j= 2n+2p

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

then what

jimthompson5910 (jim_thompson5910):

move the 2j over

OpenStudy (anonymous):

subtract it

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yes subtract 2j from both sides

OpenStudy (anonymous):

then divide both sides by 2?

jimthompson5910 (jim_thompson5910):

divide both sides by 2m

jimthompson5910 (jim_thompson5910):

to isolate k

OpenStudy (anonymous):

\[k =\frac{ n }{ m } +\frac{ p }{ m } + \frac{ j }{ m }\]

jimthompson5910 (jim_thompson5910):

it should be \[\Large k = \frac{n}{m}+\frac{p}{m}-\frac{j}{m}\] you can also write that as \[\Large k = \frac{n+p-j}{m}\]

jimthompson5910 (jim_thompson5910):

notice it's minus j over m and NOT plus j over m

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

thats thr final answer?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes either form works but I prefer the second form

OpenStudy (anonymous):

thanks can you help me with one other thing

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

How would you write the slope-intercept inequality for a graph

jimthompson5910 (jim_thompson5910):

so you're given a graph and they want the inequality?

OpenStudy (anonymous):

yes but if you have the two points (0,5) and (3,-1) and the line is dashed and shaded like this